OpenStudy (javaz):

Your fixed expenses are $1,151.65/month. You saved 4 months' worth in an emergency fund, investing 25% in a savings account at a 3.3% APR and the rest in a 60-day CD at a 4.3% APR. How much total interest accrues over 60 days?

OpenStudy (niksbhalla14mar):

4 Months Worth = 1151.65 x 4 = 4606.60 25% of 4606.60 is 1151.65 which is invested at the rate of 3.3% 75% of 4606.60 is 3454.95 which is invested at the rate of 4.3% Investment ratio is 1:3 So net interest rate would be also in the opposite ratio as per the law of alligation. So Net rate is 4.05% two months = 2/12 PRT/100 = 4606.6 x 4.05% x (2/12) = $31 Answer is $31

OpenStudy (javaz):

thank you